Hi, I'm not sure if I should be in this section, but here I am. We are studying compound inequalities etc... in my class right now. My question is a word problem that I don't even know where to start. Can someone please help me?
The body mass index, I, can be used to determine an individual's risk of heart diease. An index of 25 indicates low risk. The body mass index is given by the formula, or model, I=700W/H^2, where W= weight and H = height, in inches. Francis weighs 189 pounds and stands 68 inches tall. What is his approximate body mass? Find an inequality describing all weights W that Francis can have and be in the low risk category. Thank-you for any help.

Answer by stanbon(75887)

The body mass index, I, can be used to determine an individual's risk of heart diease. An index of 25 indicates low risk.
----------------------------------------------------------
The body mass index is given by the formula, or model, I=700W/H^2, where W= weight and H = height, in inches.
-----------------------------------------------------------
Francis weighs 189 pounds and stands 68 inches tall. What is his approximate body mass?
I = 700*189/68^2 = 28.61
-----------------------------
Find an inequality describing all weights W that Francis can have and be in the low risk category.
700*W/H^2 <= 25
700*W/68^2 <= 25
700W <= 115600
Weight <= 156.143 lbs.
=============================
Cheers,
Stan H.